OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step?

OpenStudy (ash2326):

First step would be to subtract the two equations, that will eliminate y and then we can find x easily. \[2x+y=8\] \[x+y=4\] Subtract 2nd from 1st \[2x+y-(x+y)=8-4\] We get \[x=4\] Now to find y, substitute x in any of the equations We have \[x+y=4\] x=4 so \[y=0\]

OpenStudy (anonymous):

so it would be 2x+y=8------> 2x+y=8 (x+y=4) (-1)--> -x-y=-4

OpenStudy (ash2326):

Yeah